OpenStudy (b77w):

How to find the absolute value of 1-5i

OpenStudy (anonymous):

The absolute value of a+bi is: \[|a+bi|=\sqrt{a^2 + b^2}\]So in this case, a=1 and b=-5. So plug those in and what do you get?

OpenStudy (b77w):

So the absolute value is 26?

OpenStudy (anonymous):

Don't forget your square root. The absolute value is the square root of 26.

OpenStudy (b77w):

Ok so sqrt(26) is my answer

OpenStudy (anonymous):

Right

OpenStudy (b77w):

Thanks again!

OpenStudy (anonymous):

No problem, glad I could help